OpenStudy (daniel.ohearn1):

I'm a fan of whomever can explain how to integrate (sin(2x^3)(sec(x))^2)^3 dx

OpenStudy (alekos):

Have you tried Wolfram?

OpenStudy (daniel.ohearn1):

Wolfram can't handle it, as is, yes. It just show's: Computing...

OpenStudy (freckles):

is (sec(x))^2)^3 a type-o just asking since this as it is written is just sec^6(x)?

OpenStudy (faiqraees):

Please use the equation editor here to write the question

OpenStudy (hexiflexidexi):

Can you express your equation more clearly? (Is it\[\sin(2x^{3})(\sec ^{2} x)^{3}\] ?) I guess it's either integration by parts or by trigonometric identities/substitution?

hartnn (hartnn):

i think the question would be \(\Large\int \sin^32x(\sec^2x)^3 dx\) is it?